# A Brain-to-Population Graph Learning Framework for Diagnosing Brain Disorders

### v1

@@ 5.4.2 Effectiveness of Node Reassignment and Group Size Sensitivity
As shown in Table 6, removing the node reassignment module leads to a noticeable drop in model performance, with accuracy and AUC decreasing to 70.6% and 70.4%, respectively. Replacing the adaptive reassignment mechanism with a WGAT layer yields some improvement over the removal case, indicating that incorporating structural grouping is beneficial. However, this variant still falls short of the full ANR-GAT model, which achieves the highest accuracy of 73.7% and AUC of 73.2%. These results demonstrate that our adaptive node reassignment strategy plays a critical role in enabling the model to learn more flexible and individualized brain region groupings, thereby enhancing the robustness and predictive power of the brain graphs.
@@ 5.6 Analysis of Loss Function in Two Stage
Stage 1 w/o similarity 73.4 ± 0.9 72.8 ± 1.0 81.3 ± 1.6 64.3 ± 2.8
Stage 1 w/ similarity 73.7 ± 0.6 73.2 ± 0.6 80.6 ± 1.8 65.9 ± 2.1
Stage 2 w/o similarity 82.0 ± 0.6 81.6 ± 0.6 86.6 ± 1.2 76.6 ± 1.5
Stage 2 w/ similarity 83.4 ± 0.4 83.0 ± 0.3 88.9 ± 0.5 77.1 ± 0.1
@@ 5.7.2 Group Reassignment
These region groupings emerged from the model’s learned co-assignment patterns, suggesting that they may share similar feature representations or jointly aid in distinguishing brain disorders.
@@ 6 Discussion and Conclusion
Despite these promising findings, several limitations remain. The brain region embeddings used in our method are derived from general-purpose large language models, which may not fully reflect domain-specific anatomical and functional nuances. Incorporating structured neuroscience knowledge, such as expert-curated ontologies or dedicated anatomical corpora, could yield more robust and biologically grounded representations. Additionally, the variability inherent in large language models introduces uncertainty that requires further control, especially when generalizing across datasets. Future work could explore domain-adaptive fine-tuning strategies and extend the framework to other imaging modalities or clinical conditions to further validate its utility and versatility.

### v2

@@ 5.6.2 Effectiveness of Node Reassignment and Group Size Sensitivity
As shown in Table 6, removing the node reassignment module leads to a noticeable drop in model performance, with accuracy and AUC decreasing to 70.6% and 70.4%, respectively. Replacing the adaptive reassignment mechanism with a WGAT layer yields some improvement over the removal case, indicating that incorporating structural grouping is beneficial. However, this variant still falls short of the full ANR-GAT model, which achieves the highest accuracy of 73.7% and AUC of 73.2%. These results demonstrate that our adaptive node reassignment strategy plays a critical role in learning flexible and individualized brain region groupings while dynamically updating semantic edges based on empirical functional connectivity.
@@
5.6.3 Parameter Sensitivity and Computational Trade-off
The computational cost in brain representation learning is primarily influenced by two factors: the hidden dimension of the GAT layers and the semantic similarity threshold used to construct the brain graph. These parameters jointly determine the balance between model expressiveness, graph sparsity, and computational efficiency. In what follows, we provide a detailed analysis of the threshold, which directly affects both graph structure and computational cost.
The semantic similarity threshold controls the density of the constructed brain graph. A smaller value preserves more connections, leading to higher GPU memory usage and longer training time, and may introduce noisy or redundant connections that weaken the quality of learned representations. Conversely, a larger threshold prunes more edges, reducing the computational cost but potentially discarding informative relationships among brain regions. To examine this trade-off, we conducted experiments with threshold values from 0.4 to 0.8 in the brain representation learning stage across different datasets, each using its best-performing configuration. As shown in Figure 7, moderate thresholds consistently yield the best results, capturing meaningful connectivity while maintaining computational efficiency.
Figure 7: Accuracy and AUC performance of ANR-GAT model with varying threshold values (0.4-0.8) on ABIDE, ADHD200, and Rest-meta-MDD datasets.
While the other experiment parameters are also primarily selected by model performance, the corresponding GPU utilization and training time under the adopted settings are reported to assess computational feasibility. As shown in Table 7, GPU memory usage is well within the available resources, and the per-epoch training time remains consistently reasonable across datasets.
Table 7: Stage 1 parameter settings and measured computational cost (maximum GPU utilization recorded in wandb, and training time per epoch).
Dataset Hidden channel Threshold
Max GPU
utilization (%)
Training time
/ epoch (s)
ABIDE 512 0.60 59.53 1.18–1.24
ADHD-200 256 0.50 47.98 0.85–0.89
Rest-meta-MDD 512 0.60 63.34 1.94–2.00
These observations confirm that the adopted configurations achieve a reasonable balance between model performance and computational efficiency.
@@ 5.7.2 Impact of Phenotypic Feature Fusion
5.7.3 Ablation Study on the Use of Phenotypic Data
To evaluate the contribution of phenotypic information, we conducted an ablation experiment where only neuroimaging data were used in the second stage. In this setting, population graph edges were constructed solely based on the similarity of learned brain features, and all phenotypic variables were excluded from the fusion process.
Table 12: Performance comparison with and without phenotypic data.
Method ACC (%, $\uparrow$) AUC (%, $\uparrow$) SPE (%, $\uparrow$) SEN (%, $\uparrow$)
Without Phenotypic Data $76.7\pm 0.7\%$ $76.5\pm 0.6\%$ $73.4\pm 2.2\%$ $79.5\pm 2.5\%$
With Phenotypic Data $\mathbf{83.4\pm 0.4\%}$ $\mathbf{83.0\pm 0.3\%}$ $\mathbf{77.1\pm 0.1\%}$ $\mathbf{88.9\pm 0.5\%}$
As shown in Table 12, even without phenotypic information, the two-stage framework outperforms the independently trained brain feature model. This observation is consistent with previous findings suggesting that imaging data alone provide substantial predictive power for disease classification (Shi et al., 2022). Nevertheless, incorporating non-imaging variables further enhances the model’s capacity to capture inter-individual variability and improves overall classification performance.
@@ 5.8 Analysis of Loss Function in Two Stage
Stage 1 w/o similarity 73.4 $\pm$ 0.9 72.8 $\pm$ 1.0 81.3 $\pm$ 1.6 64.3 $\pm$ 2.8
Stage 1 w/ similarity 73.7 $\pm$ 0.6 73.2 $\pm$ 0.6 80.6 $\pm$ 1.8 65.9 $\pm$ 2.1
Stage 2 w/o similarity 82.0 $\pm$ 0.6 81.6 $\pm$ 0.6 86.6 $\pm$ 1.2 76.6 $\pm$ 1.5
Stage 2 w/ similarity 83.4 $\pm$ 0.4 83.0 $\pm$ 0.3 88.9 $\pm$ 0.5 77.1 $\pm$ 0.1
@@ 5.9.2 Group Reassignment
These region groupings emerged from the model’s learned co-assignment patterns, suggesting that they may share similar feature representations or jointly aid in distinguishing disease.
@@ 6 Discussion and Conclusion
Despite these promising findings, several limitations remain. The brain region embeddings used in our method are derived from general-purpose large language models, which may not fully reflect domain-specific anatomical and functional nuances. Incorporating structured neuroscience knowledge, such as expert-curated ontologies or dedicated anatomical corpora, could yield more robust and biologically grounded representations. Additionally, the variability inherent in large language models introduces uncertainty that requires further control and systematic validation. Future work could explore domain-adaptive fine-tuning strategies, and establish comprehensive protocols to ensure the factual consistency of generated descriptions. It would also be valuable to compare different atlases, including anatomical and multi-atlas schemes, and to extend the framework to other imaging modalities or clinical conditions for a more comprehensive evaluation of its general utility. In addition, integrating advanced GNN architectures, such as contrastive and self-supervised learning, could enhance the capacity to model complex relationships and improve generalizability.
CRediT authorship contribution statement
Qianqian Liao: Conceptualization, Methodology, Formal analysis, Visualization, Writing-Original draft, Writing-Review &\& Editing. Wuque Cai: Writing-Original draft, Writing-Review &\& Editing, Methodology. Hongze Sun: Visualization, Formal analysis, Writing-Review &\& Editing. Dongze Liu: Data Curation, Formal analysis. Duo Chen: Funding acquisition, Visualization. Dezhong Yao: Conceptualization, Funding acquisition, Supervision, Writing-Review &\& Editing. Daqing Guo: Conceptualization, Supervision, Funding acquisition, Resources, Writing-Original draft, Writing-Review &\& Editing.